There are two nos 6 and 9. We have to find thier AM, GM and HM using compass and ruler only.

Grade:12

3 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please ask the question and kindly repost

Regards
Vikas TU
4 years ago
Dear student 
Question is not clear 
Please attach an image, 
We will happy to  help you 
Good Luck
Cheers
Arun
4 years ago
Dear student
 
AM = (6 +9 )/2 = 15/2
GM = √6*9 = √54 = 3 √6
HM = 2 * 6*9/(6+9) = 108/15 = 36/5
 
Hope it helps
 
In casw of any difficulty, please feel free to ask again